OpenStudy (anonymous):

Write an equation, in slope-intercept form, of the line through the pair of points: (8,-4) and (0,3)

OpenStudy (watchmath):

Slope\(=\frac{3+4}{0-8}=-\frac{7}{8}\) So the line is of the form \(y=-\frac{7}{8}+b\) From |(0,3\) we know that the y-intercept is 3. So b=3. Thus \(y=\frac{-7}{8}x+3\)

OpenStudy (anonymous):

thank you!!!

OpenStudy (anonymous):

help me